# m1: Unleash the Potential of Test-Time Scaling for Medical Reasoning with Large Language Models

### v1

@@ Abstract
Figure 1: Test-time scaling of m1 series. Each plot shows accuracy (%) vs. reasoning token budget for different m1 model variants on various medical QA datasets. All models improve steadily as the thinking length increases, with the 32B model reaching the best accuracy. The linear regression lines are marked in dot line with their 95% confidence interval.
@@ 1 Introduction
Test-time scaling has emerged as a promising direction to enhance LLM reasoning by enabling models to “think more” during inference (Yang et al., 2025). OpenAI’s o1 (Jaech et al., 2024) demonstrated that significantly extending an LLM’s chain-of-thought can yield remarkable gains in problem-solving ability in both STEM fields and the medical domain (Muennighoff et al., 2025; Xie et al., 2024), but the exact methodology was not disclosed, spurring many replication efforts. Among the most successful replication attempts is the open-source s1 method (Muennighoff et al., 2025), which achieved remarkable results through a surprisingly simple approach. By fine-tuning a 32B parameter model on just 1K carefully curated examples with reasoning traces and implementing an inference control mechanism via a ”Wait” token, s1 enabled the model to effectively double-check its work. This simple approach produced state-of-the-art results on challenging mathematical benchmarks, e.g., outperformed OpenAI’s o1-preview by up to 27%.
Despite these advances, applying test-time scaling to the medical domain remains largely underexplored (Jiang et al., 2025). The medical domain presents unique challenges for LLMs: questions often involve multi-step logical reasoning, accurate recall of medical knowledge, and careful consideration to avoid unsafe or harmful answers (Chen et al., 2024b). As the medical field fundamentally differs from mathematical tasks in terms of knowledge representation and decision-making processes, the effectiveness of test-time scaling for medical reasoning remains uncertain.
While advanced proprietary models like GPT-4 (Hurst et al., 2024) and Med-PaLM (Singhal et al., 2022) have achieved expert-level scores on medical exams (Zhang et al., 2024), open-source medical LLMs still struggle to reliably solve complex medical problems. Improving reasoning in these models is critical, as healthcare applications demand not just factual accuracy but robust diagnostic and therapeutic reasoning capabilities. Existing medical reasoning LLMs such as HuatuoGPT-o1 (Chen et al., 2024b) typically rely on computationally intensive methods like reinforcement learning with verification mechanisms. This raises a key question: Can a simple test-time scaling strategy, with minimal fine-tuning, also unlock strong medical reasoning?
@@
To better understand the impact of test-time scaling on medical reasoning in LLMs, we conduct a fine-grained study, systematically examining the effects of thinking budgets, inference techniques, data curation, and model capacity. While increasing the token budget consistently improves performance, we identify an optimal reasoning threshold of approximately 4K tokens, beyond which accuracy declines due to overthinking. In addition to increasing the token budget (Figure 1), reasoning can also be extended through budget forcing, wherein the model iteratively prolongs its thought process during inference (Muennighoff et al., 2025). However, unlike in mathematical reasoning—where iterative refinement often enhances accuracy—forcing additional reasoning in medical QA yields limited benefits and, in some cases, even degrades performance. This occurs when models with erroneous knowledge reconsider correct responses during extended reasoning, ultimately arriving at incorrect conclusions.
@@
Empirically, m1 exhibits substantial performance despite its minimalism. Our 7B model fine-tuned on 23K examples (m1-7B-23K) attains new state-of-the-art accuracy of 60.32% among in-domain and out-domain medical exam datasets, surpassing previously established specialized models of similar scale such as HuatuoGPT-o1-7B/8B (trained with complex RL on 40K instances) (Chen et al., 2024b) and UltraMedical-8B (trained on hundreds of thousands of medical instructions) (Zhang et al., 2024). Furthermore, our larger 32B model trained with only 1K fine-tuning samples (m1-32B-1K) achieves performance comparable to 2X bigger resource-exhausted models (around 70B parameters with high training costs), underscoring the efficiency of our test-time scaling approach. All data, code, and models are publicly available to encourage future exploration in optimizing inference strategies in clinical AI applications.
@@ 2 Related Works
There is a growing interest in techniques that enhance an LLM’s reasoning without altering its weights, by allocating more computation at inference time (Jaech et al., 2024; Meng et al., 2023; Hu et al., 2025). A basic form is chain-of-thought prompting, e.g. instructing the model to “think step by step,” which often improves performance on complex tasks (Wei et al., 2022). More explicit approaches include generating multiple solutions and using majority voting or self-consistency to pick an answer, or employing search-based strategies with verifiers and lookahead. These methods trade extra inference passes for accuracy gains. In contrast, sequential test-time scaling keeps a single reasoning thread but makes it longer. OpenAI’s o1 model hinted at the power of simply extending the reasoning length (Jaech et al., 2024). Muennighoff et al. (2025) formalized this by fine-tuning an LLM to utilize special “Wait” tokens, which allow controlling response length during inference. Their budget forcing method (described below) proved more effective than parallel voting strategies. Other recent research has proposed optimizing the allocation of test-time compute, for example finding an optimal stopping length per problem to avoid overthinking (Yang et al., 2025). Our work builds directly on the simple test-time scaling idea (Muennighoff et al., 2025; Aggarwal & Welleck, 2025) – we apply it to a new domain (medicine) and confirm its benefits in a very different setting. We focus on single-trace sequential reasoning, noting that it is complementary to orthogonal advances like tool use or retrieval augmentation.
@@
The success of GPT-4 in medical exams (Zhang et al., 2024) has spurred numerous open efforts to train medical domain LLMs. Early approaches centered on domain-specific pre-training: e.g. Wu et al. (2024) and Qiu et al. (2024) pre-trained Llama models on medical text corpora (MIMIC-III (Johnson et al., 2016), PubMed  [^1] https://pubmed.ncbi.nlm.nih.gov/, etc.) to inject medical knowledge. While this improves knowledge recall, the gains on reasoning-heavy tasks were limited (Jiang et al., 2025; Wu et al., 2025). More recent projects emphasize instruction tuning and reinforcement learning specialized for medicine. For example, OpenBioLLM was fine-tuned with expert-validated instructions and Direct Preference Optimization, and reportedly outperformed GPT-4 and Med-PaLM-2 on several biomedical QA benchmarks (Pal & Sankarasubbu, 2024). Med42 is another open model suite that achieved impressive results, even exceeding GPT-4.0 on many multi-choice medical QA tasks (Christophe et al., 2024). To push reasoning ability further, some works incorporate explicit reasoning supervision or verification. HuatuoGPT-o1 introduced verifiable medical problem-solving: they constructed 40K problems with known solutions and used a two-stage training (SFT + RL with a verifier) to train a 70B model (Chen et al., 2024b). This model achieved new state-of-the-art results on medical reasoning benchmarks, outperforming both general and prior medical LLMs (Zhang et al., 2024). UltraMedical built a massive dataset of  410K mixed manual/synthetic instructions for biomedicine (Jiang et al., 2025), and fine-tuned Llama-3 models with supervised and preference learning. The 70B UltraMedical model reached 86.5% accuracy on MedQA, nearly matching Med-PaLM2 (Singhal et al., 2025) and GPT-4 (Achiam et al., 2023). In contrast, our approach remains lightweight as we do not introduce new RL or verification components, and our data size (1K–23K) is tiny, yet through test-time scaling, we achieve competitive results with these state-of-the-art models. We hope this encourages more exploration of inference-time techniques as an efficient alternative for domain-specific LLMs.
@@ 3 Method
Figure 2: An overview of our data curation and training pipeline. We start with 196K raw medical QA examples, apply difficulty filtering (retaining 37K that Qwen2.5-7B-Instruct (Yang et al., 2024) or its 32B version cannot solve), then use DeepSeek-R1 (Guo et al., 2025) to generate reasoning and keep correct solutions (m23K). We perform diversity sampling to select a 1K high-quality subset (m1K). These datasets are used to fine-tune base models (Qwen2.5 7B and 32B Instruct) via Supervised Fine-Tuning (SFT), resulting in the m1 models (m1-7B-1K, m1-7B-23K, m1-32B-1K).
Our approach consists of three parts: 1) Data curation: selecting and generating a high-quality set of medical QA examples with detailed reasoning (Section 3.1), 2) Model training: Supervised Fine-Tuning (SFT) of base LLMs on this data (Section 3.2), and 3) Inference: test-time control of the model’s reasoning length (Section 3.3). Figure 2 provides an overview of the full pipeline.
@@ 3.1 Data curation
To construct the training data for m1 through a multi-step refinement process, we begin with a large pool of approximately 196K medical QA samples compiled from public datasets: MedMCQA (Pal et al., 2022), MedQA-USMLE (Jin et al., 2021), HeadQA (Vilares & Gómez-Rodríguez, 2019), and PubMedQA (Jin et al., 2019). These include multiple-choice questions from medical exams as well as open-ended research questions. All samples are decontaminated against the evaluation data in Section 4.1. More details are presented in Appendix A.
@@
Following (Muennighoff et al., 2025), we identify a subset of solvable yet non-trivial problems by performing difficulty filtering using two strong base models. Specifically, we use Qwen2.5-Instruct (Yang et al., 2024) (an open general LLM) of 7B and 32B parameters to attempt each question. We filter a question if either Qwen-7B or Qwen-32B answers it correctly. This heuristic retains questions that are challenge to solve, eliminating those that are too easy for either models. Difficulty filtering pruned the dataset from 196K down to  37K samples.
@@
We employ DeepSeek-R1 (Guo et al., 2025), a state-of-the-art open reasoning LLM, to generate a chain-of-thought and final answer for each of the 37K questions. DeepSeek-R1 was chosen for its robust reasoning capability (it’s comparable to OpenAI’s o1 (Jaech et al., 2024) in multi-step problem solving). For each question, we prompt DeepSeek-R1 to produce a detailed solution explanation ending in a definitive answer. We then apply solution validation: we only keep those instances where DeepSeek-R1’s final answer is correct (matching the ground-truth). This yields a set of  23K high-quality “thinking - answers”, where each question now paired with a verified-correct reasoning process. This step ensures our training data predominantly consists of valid reasoning, while incorrect chains are discarded. More details are presented in Appendix A and B.1.
@@
We design a diversity sampling strategy to construct a well-balanced and enriched subset for training our primary model. This process highlights two key components: domain balance and dataset balance. First, we ensure domain balance by annotating each sample with Medical Subject Headings (MeSH) categories  [^2] https://www.ncbi.nlm.nih.gov/mesh/, enabling systematic coverage across medical specialties (e.g., cardiology, neurology) and question types. Second, we address dataset imbalance through stratified sampling, first selecting domains, then source datasets, and finally individual samples (see Appendix Tables 3 and 4 for distributions). We perform stratified sampling at the dataset level to address the imbalance in sample counts across datasets (Appendix, Table 34). Specifically, we first sample a domain, then sample a dataset, and finally roll-out a sample. The process is repeated until there are 1K samples (m1K), which will be served as our core training set for m1. The remaining 23K difficult samples (m23K) can be used to augment training or for ablations. We provide summary statistics of the final data in Appendix A.
@@ 3.2 Model Training
We fine-tune three model variants, corresponding to two model sizes (7B and 32B) and two training set sizes (1K and 23K). For each, we use the pre-trained Qwen2.5-Instruct model as the initialization. Qwen2.5 is a recent high-performance open LLM (Yang et al., 2024); using it as our base ensures strong general language ability and allows us to focus on injecting medical reasoning. We format each training example in a “question $\to$ reasoning $\to$ answer” style. This format teaches the model to produce a coherent reasoning process and then give the answer. Using this data, we perform SFT for each model:
@@ 3.3 Inference
At inference time, we employ test-time scaling by managing the model’s generation of the chain-of-thought. Specifically, we define a thinking budget: a maximum number of tokens the model is allowed to generate before producing a final answer. By allocating a larger budget, we give the model more “thinking space” to potentially reason through the problem. If the model would naturally finish its reasoning early, we intervene to use the budget fully.
Besides, we apply budget forcing technique to extend the thinking process of the model: when the model outputs end-of-think token indicating the end of thinking before reaching the token budget, we replace it with “Wait.” and force the model to keep thinking generation. According to (Muennighoff et al., 2025), this method often leads the model to double-check or refine its initial answers for math problems.
@@ 4 Experiments
Table 1: Baseline Comparisons. We report accuracy (%) on each evaluation dataset for various models. Our m1 models (in bold) are shown in the $\leq 10B$ group (m1-7B variants) and $>10B$ group (m1-32B). “+CoT” indicates using chain-of-thought prompting at inference for that base model. We mark Green color within each parameter groupr: the deeper the color, the higher the accuracy. For header abbreviations, please refer to Section 4.1.
@@ 4.1 Evaluation Settings
We evaluate on nine medical QA benchmarks, grouped into In-Distribution and Out-of-Distribution tests. We measure accuracy for all datasets. 1) In-Distribution tests, the companied test splits from our training data, include MedMCQA (Pal et al., 2022) (MedMC); MedQA-USMLE (Jin et al., 2021) (MedQA), and PubMedQA (Jin et al., 2019) (PubMed). 2) Out-of-Distribution tests, which not included in training and stylistically distinct, assessing m1’s reasoning generalization: medical related questions from MMLU-Pro (Wang et al., 2024) (MMLU-P) and GPQA (Rein et al., 2024), small QA sets from Lancet and the New England Journal of Medicine (NEJM); 4 Options (MEdB (4)) and 5 Options (MedB (5)) splits from the MedBullets platform (Chen et al., 2024a); and MedXpertQA (Zuo et al., 2025) (MedX).
@@
We compare our models against a variety of general and specialized medical LLM baselines: 1) general base instruct models Qwen2.5-7B, tested both as-is and with chain-of-thought prompting (+CoT); 2) specialized medical models including MedLlama3  [^3] https://huggingface.co/johnsnowlabs/, OpenBioLLM (Pal & Sankarasubbu, 2024), MMed-Llama (Qiu et al., 2024), Med42 (Christophe et al., 2024), UltraMedical (Zhang et al., 2024); and 3) medical reasoning model HuatuoGPT-o1 (Chen et al., 2024b) undergoes complex RL training.
Additionally, we compare to state-of-the-art large open medical LLMs ( 10B), including Med42-70B (Christophe et al., 2024), OpenBioLLM-70B (Pal & Sankarasubbu, 2024), UltraMedical-70B (Zhang et al., 2024), HuatuoGPT-o1-70B/72B (Chen et al., 2024b), and baseline Qwen2.5 models (32B, 72B) with their respective +CoT versions. HuatuoGPT-o1 and UltraMedical employ complex training strategies involving reinforcement learning or expert feedback; therefore, matching or exceeding their performance with our simpler test-time scaling method underscores its effectiveness.

### v2

@@ Abstract
keywords:
Medical, Reasoning, Large Language Models, Test-Time Scaling, Health Care
Figure 1: Test-time scaling of m1 series. Each plot shows accuracy (%) vs. reasoning token budget for different m1 model variants on various medical QA datasets. All models improve steadily as the thinking length increases, with the 32B model reaching the best accuracy. The linear regression lines are dotted with 95% CIs.
Data and Code Availability
Our code, models, and data are publicly available at https://github.com/UCSC-VLAA/m1.
Institutional Review Board (IRB)
Our research does not require IRB approval.
@@ 1 Introduction
Test-time scaling has emerged as a promising direction to enhance LLM reasoning by enabling models to “think more” during inference yang2025towards. OpenAI’s o1 jaech2024openai demonstrated that significantly extending an LLM’s chain-of-thought can yield remarkable gains in problem-solving ability in both STEM fields and the medical domain muennighoff2025s1; xie2024preliminary, but the exact methodology was not disclosed, spurring many replication efforts. Among the most successful replication attempts is the open-source s1 method muennighoff2025s1, which achieved remarkable results through a surprisingly simple approach. By fine-tuning a 32B parameter model on just 1K carefully curated examples with reasoning traces and implementing an inference control mechanism via a ”Wait” token, s1 enabled the model to effectively double-check its work. This simple approach produced state-of-the-art results on challenging mathematical benchmarks, e.g., outperformed OpenAI’s o1-preview by up to 27%.
Despite these advances, applying test-time scaling to the medical domain remains largely underexplored jiang2025meds. The medical domain presents unique challenges for LLMs: questions often involve multi-step logical reasoning, accurate recall of medical knowledge, and careful consideration to avoid unsafe or harmful answers chen2024huatuogpt. As the medical field fundamentally differs from mathematical tasks in terms of knowledge representation and decision-making processes, the effectiveness of test-time scaling for medical reasoning remains uncertain.
While advanced proprietary models like GPT-4 hurst2024gpt and Med-PaLM singhal2022large have achieved expert-level scores on medical exams zhang2024ultramedical, open-source medical LLMs still struggle to reliably solve complex medical problems. Improving reasoning in these models is critical, as healthcare applications demand not just factual accuracy but robust diagnostic and therapeutic reasoning capabilities. Existing medical reasoning LLMs such as HuatuoGPT-o1 chen2024huatuogpt typically rely on computationally intensive methods like reinforcement learning with verification mechanisms. This raises a key question: Can a simple test-time scaling strategy, with minimal fine-tuning, also unlock strong medical reasoning?
@@
To better understand the impact of test-time scaling on medical reasoning in LLMs, we conduct a fine-grained study, systematically examining the effects of thinking budgets, inference techniques, data curation, and model capacity. While increasing the token budget consistently improves performance, we identify an optimal reasoning threshold of approximately 4K tokens, beyond which accuracy declines due to overthinking. In addition to increasing the token budget (Figure 1), reasoning can also be extended through budget forcing, wherein the model iteratively prolongs its thought process during inference muennighoff2025s1. However, unlike in mathematical reasoning—where iterative refinement often enhances accuracy—forcing additional reasoning in medical QA yields limited benefits and, in some cases, even degrades performance. This occurs when models with erroneous knowledge reconsider correct responses during extended reasoning, ultimately arriving at incorrect conclusions.
@@
Our 7B model fine-tuned on 23K examples (m1-7B-23K) attains new state-of-the-art accuracy of 60.32% among in-domain and out-domain medical exam datasets, surpassing previously established specialized models of similar scale such as HuatuoGPT-o1-7B/8B (trained with complex RL on 40K instances) chen2024huatuogpt and UltraMedical-8B (trained on hundreds of thousands of medical instructions) zhang2024ultramedical. Furthermore, our larger 32B model trained with only 1K fine-tuning samples (m1-32B-1K) achieves performance comparable to 2X bigger resource-exhausted models (around 70B parameters with high training costs), underscoring the efficiency of our test-time scaling approach. All data, code, and models are publicly available to encourage future exploration in optimizing inference strategies in clinical AI applications.
@@ 2 Related Works
There is a growing interest in techniques that enhance an LLM’s reasoning without altering its weights, by allocating more computation at inference time jaech2024openai; meng2023deepscaler; OpenReasonerZero2025. A basic form is chain-of-thought prompting, e.g. instructing the model to “think step by step,” which often improves performance on complex tasks wei2022chain. More explicit approaches include generating multiple solutions and using majority voting or self-consistency to pick an answer, or employing search-based strategies with verifiers and lookahead. These methods trade extra inference passes for accuracy gains. In contrast, sequential test-time scaling keeps a single reasoning thread but makes it longer. OpenAI’s o1 model hinted at the power of simply extending the reasoning length jaech2024openai. muennighoff2025s1 formalized this by fine-tuning an LLM to utilize special “Wait” tokens, which allow controlling response length during inference. Their budget forcing method (described below) proved more effective than parallel voting strategies. Other recent research has proposed optimizing the allocation of test-time compute, for example finding an optimal stopping length per problem to avoid overthinking yang2025towards. Our work builds directly on the simple test-time scaling idea muennighoff2025s1; aggarwal2025l1 by extending thinking traces with “wait” — we apply it to a new domain (medicine) and confirm its benefits in a very different setting. We focus on single-trace sequential reasoning, noting that it is complementary to orthogonal advances like tool use or retrieval augmentation.
@@
The success of GPT-4 in medical exams zhang2024ultramedical has spurred numerous open efforts to train medical domain LLMs lucas2024reasoning; xiong2024improving. Early approaches centered on domain-specific pre-training: e.g. wu2024pmc and qiu2024towards pre-trained Llama models on medical text corpora (MIMIC-III johnson2016mimic, PubMed  [^1] https://pubmed.ncbi.nlm.nih.gov/, etc.) to inject medical knowledge. While this improves knowledge recall, the gains on reasoning-heavy tasks were limited jiang2025meds; medreason. More recent projects emphasize instruction tuning and reinforcement learning specialized for medicine. For example, OpenBioLLM was fine-tuned with expert-validated instructions and Direct Preference Optimization, and reportedly outperformed GPT-4 and Med-PaLM-2 on several biomedical QA benchmarks pal2024openbiollms. Med42 is another open model suite that achieved impressive results, even exceeding GPT-4.0 on many multi-choice medical QA tasks christophe2024med42. To push reasoning ability further, some works incorporate explicit reasoning supervision or verification. HuatuoGPT-o1 introduced verifiable medical problem-solving: they constructed 40K problems with known solutions and used a two-stage training (SFT + RL with a verifier) to train a 70B model chen2024huatuogpt. This model achieved new state-of-the-art results on medical reasoning benchmarks, outperforming both general and prior medical LLMs zhang2024ultramedical. UltraMedical built a massive dataset of  410K mixed manual/synthetic instructions for biomedicine jiang2025meds, and fine-tuned Llama-3 models with supervised and preference learning. The 70B UltraMedical model reached 86.5% accuracy on MedQA, nearly matching Med-PaLM2 singhal2025toward and GPT-4 achiam2023gpt. In contrast, our approach remains lightweight as we do not introduce new RL or verification components, and our dataset size (1K-23K) is relatively small, yet through test-time scaling, we achieve competitive results with these state-of-the-art models. We hope this encourages more exploration of inference-time techniques as an efficient alternative for domain-specific LLMs.
@@ 3 Method
Figure 2: An overview of our data curation and training pipeline. We start with 196K raw medical QA examples, apply difficulty filtering (retaining 37K that Qwen2.5-7B-Instruct yang2024qwen2 or its 32B version cannot solve), then use DeepSeek-R1 guo2025deepseek to generate reasoning and keep correct solutions (m23K). We perform diversity sampling to select a 1K high-quality subset (m1K). These datasets are used to fine-tune base models (Qwen2.5 7B and 32B Instruct) via Supervised Fine-Tuning (SFT), resulting in the m1 models (m1-7B-1K, m1-7B-23K, m1-32B-1K).
@@ 3.1 Data Curation
To construct the training data for m1 through a multi-step refinement process, we begin with a large pool of approximately 196K medical QA samples compiled from public datasets: MedMCQA pal2022medmcqa, MedQA-USMLE jin2021disease, HeadQA vilares2019head, and PubMedQA jin2019pubmedqa. These include multiple-choice questions from medical exams as well as open-ended research questions. All samples are decontaminated against the evaluation data in Section 4.1. More details are presented in Appendix A.
@@
Following s1 muennighoff2025s1, we identify a subset of solvable yet non-trivial problems by performing difficulty filtering using two strong base models. Specifically, we use Qwen2.5-Instruct yang2024qwen2 (an open general LLM) of 7B and 32B parameters to attempt each question. We filter a question if either Qwen-7B or Qwen-32B answers it correctly. This heuristic retains questions that are challenging to solve, eliminating those that are too easy for either models. Difficulty filtering pruned the dataset from 196K down to  37K samples.
@@
We employ DeepSeek-R1 guo2025deepseek, a state-of-the-art open reasoning LLM, to generate a chain-of-thought and final answer for each of the 37K questions. DeepSeek-R1 was chosen for its robust reasoning capability (it’s comparable to OpenAI’s o1 jaech2024openai in multi-step problem solving). For each question, we prompt DeepSeek-R1 to produce a detailed solution explanation ending in a definitive answer. We then apply solution validation: we only keep those instances where DeepSeek-R1’s final answer is correct (matching the ground-truth). This yields a set of  23K high-quality “thinking - answers”, where each question now paired with a verified-correct reasoning process. This step ensures our training data predominantly consists of valid reasoning, while incorrect chains are discarded. More details are presented in Appendix A and B.1.
@@
We design a diversity sampling strategy to construct a well-balanced and enriched subset for training our primary model. This process highlights two key components: domain balance and dataset balance. First, we ensure domain balance by annotating each sample with Medical Subject Headings (MeSH) categories [^2] https://www.ncbi.nlm.nih.gov/mesh/ (See Appendix for the details), enabling systematic coverage across medical specialties (e.g., cardiology, neurology) and question types. Second, we address dataset imbalance through stratified sampling, first selecting domains, then source datasets, and finally individual samples (see Appendix Tables 3 and 4 for distributions). We perform stratified sampling at the dataset level to address the imbalance in sample counts across datasets (Appendix, Table 34). Specifically, we first sample a domain, then sample a dataset, and finally roll-out a sample. The process is repeated until there are 1K samples (m1K), which will be served as our core training set for m1. The remaining 23K difficult samples (m23K) can be used to augment training or for ablations. We provide summary statistics of the final data in Appendix A.
@@ 3.2 Model Training
We fine-tune three model variants, corresponding to two model sizes (7B and 32B) and two training set sizes (1K and 23K). For each, we use the pre-trained Qwen2.5-Instruct model as the initialization. Qwen2.5 is a recent high-performance open LLM yang2024qwen2; using it as our base ensures strong general language ability and allows us to focus on injecting medical reasoning. We format each training example in a “question $\to$ reasoning $\to$ answer” style. This format teaches the model to produce a coherent reasoning process and then give the answer. Using this data, we perform SFT for each model:
@@ 3.3 Inference
At inference time, we employ test-time scaling by managing the model’s generation of the chain-of-thought. Specifically, we define a thinking budget: a maximum number of tokens the model is allowed to generate before producing a final answer. By allocating a larger budget, we give the model more “thinking space” to potentially reason through the problem. If the model would naturally finish its reasoning early, we intervene to use the budget fully. We also apply budget forcing technique to extend the thinking process of the model: when the model outputs end-of-think token indicating the end of thinking before reaching the token budget, we replace it with “Wait.” and force the model to keep the generation of the thinking traces. According to muennighoff2025s1, this method often leads the model to double-check or refine its initial answers for math problems.
@@ 4 Experiments
Table 1: Baseline Comparisons. We report accuracy (%) on each evaluation dataset for various models. Our m1 models (in bold) are shown in the $\leq 10B$ group (m1-7B variants) and >10B>10B group (m1-32B). “+CoT” indicates using chain-of-thought prompting at inference for that base model. We mark Green color within each parameter group: the deeper the color, the higher the accuracy. For header abbreviations, please refer to Section 4.1.
@@ 4.1 Evaluation Settings
We evaluate on nine medical QA benchmarks, grouped into In-Distribution and Out-of-Distribution tests. We measure accuracy for all datasets. 1) In-Distribution tests: The accompanying test splits from our training data, include MedMCQA pal2022medmcqa (MedMC); MedQA-USMLE jin2021disease (MedQA), and PubMedQA jin2019pubmedqa (PubMed). 2) Out-of-Distribution tests: The datasets are not included in training and stylistically distinct, assessing m1’s reasoning generalization: medical related questions from MMLU-Pro wang2024mmlu (MMLU-P) and GPQA rein2024gpqa, small QA sets from Lancet and the New England Journal of Medicine (NEJM); 4 Options (MEdB (4)) and 5 Options (MedB (5)) splits from the MedBullets platform chen2024benchmarking; and MedXpertQA zuo2025medxpertqa (MedX).
@@
We compare our models against a variety of general and specialized medical LLM baselines: 1) general base instruct models Qwen2.5-7B, tested both as-is and with chain-of-thought prompting (+CoT); 2) specialized medical models including MedLlama3  [^3] https://huggingface.co/johnsnowlabs/, OpenBioLLM pal2024openbiollms, MMed-Llama qiu2024towards, Med42 christophe2024med42, UltraMedical zhang2024ultramedical; 3) medical reasoning model HuatuoGPT-o1 chen2024huatuogpt, which undergoes complex RL training.
Additionally, we compare to state-of-the-art large open medical LLMs (>10B), including Med42-70B christophe2024med42, OpenBioLLM-70B pal2024openbiollms, UltraMedical-70B zhang2024ultramedical, HuatuoGPT-o1-70B/72B chen2024huatuogpt, and baseline Qwen2.5 models (32B, 72B) with their respective +CoT versions. HuatuoGPT-o1 and UltraMedical employ complex training strategies involving reinforcement learning or expert feedback; therefore, matching or exceeding their performance with our simpler test-time scaling method underscores its effectiveness.